# TabSTAR: A Foundation Tabular Model With Semantically Target-Aware Representations

### v1

@@ Abstract
While deep learning has achieved remarkable success across many domains, it has historically underperformed on tabular learning tasks, which remain dominated by gradient boosting decision trees (GBDTs). However, recent advancements are paving the way for Tabular Foundation Models, which can leverage real-world knowledge and generalize across diverse datasets, particularly when the data contains free-text. Although incorporating language model capabilities into tabular tasks has been explored, most existing methods utilize static, target-agnostic textual representations, limiting their effectiveness. We introduce TabSTAR: a Foundation Tabular Model with Semantically Target-Aware Representations. TabSTAR is designed to enable transfer learning on tabular data with textual features, with an architecture free of dataset-specific parameters. It unfreezes a pretrained text encoder and takes as input target tokens, which provide the model with the context needed to learn task-specific embeddings. TabSTAR achieves state-of-the-art performance for both medium- and large-sized datasets across known benchmarks of classification tasks with text features, and its pretraining phase exhibits scaling laws in the number of datasets, offering a pathway for further performance improvements.
@@ 1 Introduction
In this paper, we introduce TabSTAR: a novel Tabular Foundation Model with Semantically Target-Aware Representations, [^2] Code is available at https://github.com/alanarazi7/TabSTAR. designed explicitly for end-to-end handling of purely textual features. By integrating an unfrozen text encoder at its core, TabSTAR can optimize free-text feature representations, demonstrating their clear superiority over alternative frozen embedding approaches. Additionally, it introduces a novel approach of target-aware tokens, which inject semantic information about the target variable as part of the input, allowing for efficient parameter sharing and resulting in an architecture with no dataset-specific parameters (see Figure 1). TabSTAR’s training is highly efficient [^3] Pretraining within 48 hours on a single A40 GPU. Finetuning with PEFT for low memory footprint. and its performance steadily improves with more pretraining data. Empirically, TabSTAR achieves state-of-the-art (SOTA) performance on classification benchmarks containing substantial textual content, demonstrating significant advances over GBDTs and leading TFMs.
@@ 2 Related Work
This section reviews prior work in five areas relevant to our approach. We begin with deep learning methods tailored for tabular data, which were applied to a single dataset. We then discuss cross-dataset transfer learning techniques that improve generalization by leveraging related datasets. Next, we cover the field of TFMs, which aim to generalize across diverse tasks and datasets through large-scale pretraining. Finally, we review recent work on applying large language models (LLMs) to tabular data and elaborate on existing AutoML [33] multimodal solutions.
@@
Several architectures have been proposed to enhance deep learning for tabular data [67, 44, 78, 81]. TabNet [3] and TabTransformer [42] introduced attention mechanisms into tabular deep learning, while FT-Transformer [25] and its improvement [26] jointly integrated numerical and categorical features into a transformer [76]. Other novel approaches leveraged inter-example information at inference time, with SAINT [66] proposing row-level attention between examples, Non-Parametric Transformers [47] processing the entire dataset, including labels, in a single forward pass, and TabR [27] combining a k-nearest-neighbor mechanism with a traditional Multi-Layer Perceptron (MLP) architecture. Recent works such as TabM [28] and RealMLP [38] focused on refining MLPs without an attention component. Despite these innovations, single-dataset deep learning models have not yet convincingly outperformed GBDTs [62, 30, 61]. Furthermore, none of them addressed the challenge of modeling tabular datasets with rich textual features.
@@
Deep learning was proven to shine when performing transfer learning in many machine learning domains [87]. Motivated by this success, [50, 85] proved that cross-dataset learning can boost single-dataset performance, but were limited to strict requirements such as partial overlap of feature names. To address this limitation, TransTab [79] integrated semantic understanding into feature tokenization, and XTab [86] pretrained a transformer backbone with dataset-specific parameters, proving that pretraining contributes to a stronger initialization for a downstream task. Despite their small scale, these studies demonstrated cross-dataset transfer learning’s potential, laying essential groundwork for the rise of TFMs.
@@
TFMs represent an emerging paradigm in tabular learning. While the definition is still evolving, we adopt the framing proposed by [74], which identifies key desired characteristics of TFMs: large-scale pretraining with adaptability to downstream tasks, mixed-type column support, cross-domain generalization, use of textual metadata, [^4] Contextual information such as the dataset description, column names and category names. and column-order invariance.
TabPFN [35] is recognized as the first TFM, and its successor TabPFN-v2 [37] currently sets the SOTA in tabular learning, becoming a popular approach for TFMs [58, 20]. TabPFN-v2 was the first model to consistently outperform GBDTs on medium-sized datasets, by pretraining Bayesian Prior-Data Fitted Networks (PFNs) [55] on 130 million synthetic datasets. Using ICL at inference time, it accepts up to 10,000 examples as input and predicts without updating its weights. Nonetheless, similarly to GBDTs, TabPFN-v2 uses off-the-shelf embeddings for text features, limiting its effectiveness.
CM2 [85], CARTE [45] and TP-BERTa [80] represent a shift toward semantic tabular modeling, leveraging textual signals and external knowledge at a greater scale. Unlike prior methods, these models transfer knowledge via language representations. CM2 pretrained on over 2,000 datasets, but did not focus on free-text features and used static word embeddings without further finetuning them. CARTE encodes tables as star-shaped graphs, jointly representing features by their names and values, and applies attention over the graph to capture contextual relations. While effective for high-cardinality features, it lacks exposure to longer free-text fields during pretraining and was proven useful mainly for small datasets. TP-BERTa adapts RoBERTa [51] with intra-feature attention and a tokenization scheme that maps numerical values into discrete relative-magnitude bins, to address the weakness of language models when tokenizing numbers [72]. Although it performs well, its use of dataset-specific output layers limits scalability and complicates multi-task learning. Consequently, they trained two separate models, [^5] One for classification and one for regression. A joint model for both tasks performed significantly worse. wasting potential for better cross-dataset learning. Notably, none of these approaches finetune semantic representations during downstream task training. In our work, we demonstrate that this is critical to align textual and tabular features.
@@
The remarkable success of LLMs is unprecedented [12, 56]. During the past years, several research attempts have tried to combine LLMs and tabular data. One line of work is on using LLMs directly for tabular prediction, by converting tabular data into serialized text. TabLLM [34] assessed LLMs under few-shot scenarios, while Tabula-8b [23] finetuned the Llama 3-8B model extensively on tabular data. Although useful for few-shot learning, these models are computationally expensive, [^6] Llama 3-8b has orders of magnitude more parameters than TP-BERTa, which has roughly 110M parameters. suboptimal for numerical features [72, 74], and potentially compromised on widely-used benchmarks due to prior exposure during training [8]. While current generations of LLMs weren’t adopted for tabular learning, their emergent knowledge from their pretraining could be crucial when textual features are present [74, 19]. Additionally, LLMs can be used in multiple aspects of tabular learning, as they seem to be promising synthetic data generators [9, 65], useful data cleaners [6] and clever feature engineers [36].
@@ 3 TabSTAR
In this section, we introduce TabSTAR: a Tabular Foundation Model with Semantically Target-Aware Representations. Our training framework consists of two stages: (1) Pretraining, where it is pretrained over a corpus of tabular datasets [^7] Ranging from metadata-rich, text-heavy datasets to numeric-only tables lacking column names. in a multi-task regime, mixing classification with regression tasks, then (2) Finetuning, where the pretrained model is further trained with LoRA [41] on a single downstream task. TabSTAR is designed to enable effective cross-dataset learning by applying supervised learning on the target variable in both stages. At its core, is uses an unfrozen encoder-only language model, which can invoke world knowledge acquired during the language model pretraining. [^8] Note that the language-model pretraining occurs before TabSTAR’s pretraining. Unless specified differently, the term pretraining refers to TabSTAR’s pretraining, which assumes the use of a pretrained language model. The encoder is combined with a tabular-specific architecture tailored to structured data, mitigating the known limitations of language models in tabular settings [72, 74].
TabSTAR’s architecture comprises five core modules: (1) Verbalization, mapping every feature into a textual representation composed of both the column name and value, with a special treatment to numerical features for full numerical precision; (2) Encoding, transforming semantic and numerical data into meaningful embeddings of the same dimension; (3) Fusion, integrating textual and numerical representations; (4) Interaction, modeling dependencies and relationships between different elements through self-attention and cross-element interactions; and (5) Prediction, where Interaction’s outputs are projected into a real value for regression or a probability distribution for classification. Figure 1 illustrates the architecture while Appendix A elaborates, and Appendix B discusses the training.
@@
Department Cardiology “Department: Cardiology” 0-
Report Mild chest discomfort. “Report: Mild chest discomfort.” 0-
Decision Hospitalized “Target. Decision: Hospitalized” 0-
Decision Released “Target. Decision: Released” 0-
@@
All the features and each of the target values are processed into a sequence of elements. Numerical features are processed into two inputs: a numerical one and a semantic one. The numerical input is standardized using z-scores, with outlier clipping at ±3 standard deviations. In addition, they are verbalized using quantile-based binning into 10 bins, a novel approach to mitigate the precision loss inherent in language models [72]. Appendix A.1 shows a precise example and §6 discusses different verbalization strategies. In contrast, semantic features are directly verbalized by concatenating the feature name and textual value, without any numerical representation. The target variable is also included as part of the input: In classification tasks, each of the $C$ possible values is represented by an element, constant for every example, while the true value remains hidden. For regression tasks, a single element is verbalized, carrying only the target name. Table 1 shows a toy dataset of patient records and outcomes and Table 2 shows the verbalization for the first patient.
@@
We employ a pretrained e5-small-v2 [77] embedding model for semantic encoding, chosen for its strong performance on the MTEB benchmark [54] with a relatively modest parameter count. By unfreezing half of its layers, the representations are optimized for predicting the target variable, which leads to a significant impact on TabSTAR performance (see §6). Each verbalization element is encoded independently into a semantic representation, with attention applied between tokens within each sequence element. In parallel, we encode standardized numerical values by projecting them into the same dimension using a small MLP. For the patient in Table 2, this results in a numerical embedding for Age alongside semantic representations for each of the five verbalizations.
@@
To obtain a unified representation for each sequence element, we apply a fusion block consisting of a single encoder-only transformer layer. For each numerical feature, the block attends over its numerical and semantic embeddings, producing a fused representation. In our running example, the representation of Age now jointly captures both its semantic context (the fact that the value represents age) as well as its numerical value (the patient’s age, 45, or 0.27 after standardization).
@@
The fused, semantically-rich and numerically-grounded representations of all elements interact via a 6-layer Transformer encoder [76]. Each input element is now a token, with feature tokens and target tokens all attending to each other. Unlike standard language models which integrate positional encoding, the Interaction module’s inputs are order-invariant, a desiredatum for TFMs, as defined by [74]. The encoder produces contextualized representations for each target value. In our example, this yields dedicated embeddings for the Release and Hospitalization target values. The role of these representations is to carry information about the likelihood of each value to be the true value.
@@ 4 Experiments
While TabSTAR could be pretrained on a massive scale, for this work we limit ourselves to a modest pretraining corpus focusing on classification, as we believe that TabSTAR’s inductive biases are best suited to shine in this task. We manually curate a pretraining corpus of 350 high-quality tabular datasets (253 classification, 97 regression), in a tedious process in which we uncover numerous duplications in the most popular tabular repositories, OpenML [75] and Kaggle, [^10] https://www.kaggle.com/datasets as elaborated by [73]. Our datasets are sourced from popular benchmarks [24, 45, 22, 21, 30, 60, 52, 31], which have almost non-existing representation of textual tabular datasets. Thus, we furthermore increase our corpus, focusing on classification datasets with rich semantic content. See Appendix C for more details.
@@
Tabular datasets with free-text have seen little prior research, and accordingly, benchmarks are incredibly rare. Therefore, we consider all the possible datasets from AutoML Multimodal Benchmark [60], from the analysis about free-text and high-cardinality features by [31] and from the CARTE paper [45]. After a deduplication process we end up with 50 datasets. However, there are two important limitations: first, the benchmark is heavily biased towards regression, with 36 datasets in total. Secondly, 29 out of these 36 datasets were solely contributed by the CARTE benchmark, which focuses more heavily on high-cardinality features as it was pretrained over knowledge graphs. While our main motivation is classification tasks with textual features, we decide nevertheless to evaluate on the full set of 50 datasets although it heavily biases towards regression problems and high-cardinality features, rather than classification and free-text (see Appendix D).
@@
We compare TabSTAR against a diverse set of baselines. For tree-based methods, we evaluate CatBoost [57], XGBoost (XGB) [14], and Random Forest (RF) [11] with the default configuration proposed by [25]. For CatBoost and XGBoost we consider a tuned version, where hyperparameters are optimized separately for each task using random search with 5-fold cross-validation under a 4-hour budget on 8 CPU cores. Among TFMs, we evaluate TabPFN-v2 [37] and CARTE [45]. For CARTE, we tune only the learning rate for each task, following the original paper. Since the public TabPFN-v2 model does not support text, we use their closed-sourced API client. [^11] https://github.com/PriorLabs/tabpfn-client For models lacking native support for textual features, we embed text using e5-small-v2 [77], allowing a fair comparison. For more details about the hyperparameters for each baseline as well as exclusion of models such as TP-BERTa due to potential leakage concerns, see Appendix E.
@@
Each of the 50 datasets in the benchmarks is evaluated with 10 random train-test splits (90% training, 10% testing), resulting in 500 runs per model. While 30 of the datasets have more than 10,000 examples, the evaluated TFMs have strict limitations. TabPFN-v2 employs ICL and thus receives as input at most 10,000 examples. Although CARTE imposes no size cap, [^12] In their own paper, CARTE was evaluated only over up to 2,048 examples, without scaling guarantees. it suffers from inefficiency and no preset configuration, requiring six learning-rate trials and totalling 6,000 slow GPU runs. Because of these important limitations, we consider two experiment conditions: (1) 10K: Each model is trained [^13] While TabPFN-v2 isn’t technically trained, we adopt this term for conciseness. over at most 10,000 training examples, and (2) Unlimited: We add TabSTAR-Unlimit, CatBoost-Tuned-Unlimit, and XGBoost-Tuned-Unlimit and evaluate them on the full version of the 30 datasets, [^14] We technically cap the amount of examples to 100,000 for computational efficiency. while retaining 10K baselines as weaker reference points. We exclude the untuned GBDTs and keep the same number of models as in the 10K condition.
@@
To maximize the value of cross-dataset learning, instead of pretraining TabSTAR once, we create five dataset splits. Each variant is pretrained on the 350 pretraining datasets and 40 of the benchmark datasets, while the other 10 serve exclusively as its test set. Crucially, the whole collection was carefully curated to prevent any data leakage from duplicate or overly similar datasets. For finetuning, while dataset-specific hyperparameter tuning can boost performance, we believe that robust out-of-the-box defaults are essential for TFMs and their evaluation, following Tab-PFN-v2’s approach. Therefore, we use a default hyperparameters configuration that was found robust over a disjoint set of tabular datasets, as detailed in Appendix B.2.

### v2

# TabSTAR: A Tabular Foundation Model for Tabular Data with Text Fields

@@ Abstract
While deep learning has achieved remarkable success across many domains, it has historically underperformed on tabular learning tasks, which remain dominated by gradient boosting decision trees. However, recent advancements are paving the way for Tabular Foundation Models, which can leverage real-world knowledge and generalize across diverse datasets, particularly when the data contains free-text. Although incorporating language model capabilities into tabular tasks has been explored, most existing methods utilize static, target-agnostic textual representations, limiting their effectiveness. We introduce TabSTAR: a Tabular Foundation Model with Semantically Target-Aware Representations. TabSTAR is designed to enable transfer learning on tabular data with textual features, with an architecture free of dataset-specific parameters. It unfreezes a pretrained text encoder and takes as input target tokens, which provide the model with the context needed to learn task-specific embeddings. TabSTAR achieves state-of-the-art performance for both medium- and large-sized datasets across known benchmarks of classification tasks with text features, and its pretraining phase exhibits scaling laws in the number of datasets, offering a pathway for further performance improvements. [^1] Code is available at https://github.com/alanarazi7/TabSTAR.
@@ 1 Introduction
In this paper, we introduce TabSTAR: a novel Tabular Foundation Model with Semantically Target-Aware Representations, designed explicitly for end-to-end handling of purely textual features. By integrating an unfrozen text encoder at its core, TabSTAR can optimize free-text feature representations, demonstrating their clear superiority over alternative frozen embedding approaches. Additionally, it introduces a novel approach of target-aware tokens, which inject semantic information about the target variable as part of the input, allowing for efficient parameter sharing and resulting in an architecture with no dataset-specific parameters (see Figure 1). TabSTAR’s training is highly efficient [^3] Pretraining within 48 hours on a single A40 GPU. Finetuning with PEFT for a low memory footprint. and its performance steadily improves with more pretraining data. Empirically, TabSTAR achieves state-of-the-art (SOTA) performance on classification datasets containing textual features, surpassing leading TFMs as well as GBDTs tuned for 4 hours.
@@ 2 Related Work
This section reviews prior work in supervised tabular learning. We begin with deep learning methods tailored for tabular data, which were applied to a single dataset. We then discuss cross-dataset transfer learning techniques that improve generalization by leveraging related datasets. Next, we cover the field of TFMs, which aim to generalize across diverse tasks and datasets through large-scale pretraining. Finally, we review recent work on applying large language models (LLMs) to tabular data and elaborate on existing AutoML [34] multimodal solutions. As we focus on supervised learning, we do not cover self-supervised methods [82, 6] and their downstream applications.
@@
Several architectures have been proposed to enhance deep learning for tabular data [70, 45, 81, 85]. TabNet [3] and TabTransformer [43] introduced attention mechanisms into tabular deep learning, while FT-Transformer [26] and its improvement [27] jointly integrated numerical and categorical features into a transformer [79]. Other novel approaches leveraged inter-example information at inference time, with SAINT [69] proposing row-level attention between examples, Non-Parametric Transformers [49] processing the entire dataset, including labels, in a single forward pass, and TabR [28] combining a k-nearest-neighbor mechanism with a traditional Multi-Layer Perceptron (MLP) architecture. Recent works such as TabM [29] and RealMLP [39] focused on refining MLPs without an attention component. Despite these innovations, single‐dataset deep learning models have not yet convincingly outperformed GBDTs [65, 31, 64]. Furthermore, none of them addressed the challenge of modeling tabular data with rich textual features.
@@
Deep learning has been proven to shine when performing transfer learning in many machine learning domains [91]. Motivated by this success, [52, 89] proved that cross-dataset learning can boost single-dataset performance, but were limited to strict requirements such as partial overlap of feature names. To address this limitation, TransTab [83] integrated semantic understanding into feature tokenization, and XTab [90] pretrained a transformer backbone with dataset-specific parameters, proving that pretraining contributes to a stronger initialization for a downstream task. Despite their small scale, these studies demonstrated cross-dataset transfer learning’s potential, laying essential groundwork for the rise of TFMs.
@@
TFMs represent an emerging paradigm in tabular learning. While the definition is still evolving, we adopt the framing proposed by [77], which identifies key desired characteristics of TFMs: large-scale pretraining with adaptability to downstream tasks, mixed-type column support, cross-domain generalization, use of textual metadata, [^4] Contextual information such as the dataset description, column names, and category names. and column-order invariance.
TabPFN [36] is recognized as the first TFM, and its successor TabPFN-v2 [38] currently sets the SOTA in tabular learning, becoming a popular approach for TFMs [61, 21, 54]. TabPFN-v2 was the first model to consistently outperform GBDTs on medium-sized datasets, by pretraining Bayesian Prior-Data Fitted Networks (PFNs) [58] on 130 million synthetic datasets. Using ICL at inference time, it accepts up to 10,000 examples as input and predicts without updating its weights. TabPFN-v2 inspired TabICL [61] which improved scalability, and TabDPT [54], trained on real data. However, they all rely on off-the-shelf embeddings for text features like GBDTs, limiting their performance.
CM2 [86], CARTE [47], and TP-BERTa [84] represent a shift toward semantic tabular modeling, leveraging textual signals and external knowledge at a greater scale. Unlike prior methods, these models transfer knowledge via language representations. CM2 was pretrained on over 2,000 datasets, but did not focus on free-text features and used static word embeddings without further finetuning them. CARTE encodes tables as star-shaped graphs, jointly representing features by their names and values, and applies attention over the graph to capture contextual relations. While effective for high-cardinality features, it lacks exposure to longer free-text fields during pretraining and was proven useful mainly for small datasets. TP-BERTa adapts RoBERTa [53] with intra-feature attention and a tokenization scheme that maps numerical values into discrete relative-magnitude bins, to address the weakness of language models when tokenizing numbers [75]. Although it performs well, its use of dataset-specific output layers limits scalability and complicates multi-task learning. Consequently, they trained two separate models, [^5] One for classification and one for regression. A joint model for both tasks performed significantly worse. wasting potential for better cross-dataset learning. Notably, none of these approaches finetune semantic representations during downstream task training. In our work, we demonstrate that this is critical to align textual and tabular features.
@@
The remarkable success of LLMs is unprecedented [13, 59]. During the past years, several research attempts have tried to combine LLMs and tabular data. One line of work focuses on using LLMs directly for tabular prediction by converting tabular data into serialized text. TabLLM [35] assessed LLMs under few-shot scenarios, while Tabula-8b [24] finetuned the Llama 3-8B model extensively on tabular data. Although useful for few-shot learning, these models are computationally expensive, [^6] Llama 3-8b has orders of magnitude more parameters than TP-BERTa, which has roughly 110M parameters. suboptimal for numerical features [75, 77], and potentially compromised on widely-used benchmarks due to prior exposure during training [9]. While current generations of LLMs weren’t adopted for tabular learning, their emergent knowledge from their pretraining could be crucial when textual features are present [77, 20]. Additionally, LLMs can be used in multiple aspects of tabular learning, as they seem to be promising synthetic data generators [10, 68], useful data cleaners [7], and clever feature engineers [37].
@@ 3 TabSTAR
In this section, we introduce TabSTAR: a Tabular Foundation Model with Semantically Target-Aware Representations. Our training framework consists of two stages: (1) Pretraining, where the model is pretrained over a corpus of tabular datasets [^7] Ranging from metadata-rich, text-heavy datasets to numeric-only tables lacking column names. in a multi-task regime, mixing classification with regression tasks, then (2) Finetuning, where the pretrained model is further trained with LoRA [42] on a single downstream task. TabSTAR is designed to enable effective cross-dataset learning by applying supervised learning on the target variable in both stages. At its core, it uses an unfrozen encoder-only language model, which can potentially invoke world knowledge acquired during the language model pretraining. [^8] Note that the language-model pretraining occurs before TabSTAR’s pretraining. Unless specified differently, the term pretraining refers to TabSTAR’s pretraining, which assumes the use of a pretrained language model. The encoder is combined with a tabular-specific architecture tailored to structured data, mitigating the known limitations of language models in tabular settings [75, 77].
TabSTAR’s architecture comprises five core modules: (1) Verbalization, mapping every feature into a textual representation composed of both the column name and value, with a special treatment to numerical features for full numerical precision; (2) Encoding, transforming semantic and numerical inputs into meaningful embeddings of the same dimension; (3) Fusion, integrating textual and numerical representations on each feature independently; (4) Interaction, modeling dependencies and relationships between different features through cross-feature self-attention; and (5) Prediction, where outputs are projected into a real value for regression or a probability distribution for classification. Figure 1 illustrates the architecture, Appendix A elaborates it, and Appendix B discusses the training.
@@
Department Cardiology “Department: Cardiology”
Report Mild chest discomfort. “Report: Mild chest discomfort.”
Decision Hospitalized “Target. Decision: Hospitalized”
Decision Released “Target. Decision: Released”
@@
All the features and each of the target values are processed into a sequence of elements. Numerical features are processed into two inputs: a numerical one and a semantic one. The numerical input is standardized using z-scores, with outlier clipping at ±3 standard deviations. In addition, they are verbalized using quantile discretization into 10 bins, a novel approach to mitigate the precision loss inherent in language models [75]. Appendix A.1 shows a precise example and §6 discusses different verbalization strategies. In contrast, semantic features are directly verbalized by concatenating the feature name and textual value, without any numerical representation. The target variable is also included as part of the input: In classification tasks, each of the $C$ possible values is represented by an element, constant for every example, while the true value remains hidden. For regression tasks, a single element is verbalized, carrying only the target name. Table 1 shows a toy dataset of patient records and outcomes and Table 2 shows the verbalization for the first patient.
@@
We employ a pretrained e5-small-v2 [80] embedding model for semantic encoding, chosen for its strong performance on the MTEB benchmark [57] with a relatively modest parameter count. By unfreezing the upper half of its layers, the representations are optimized for predicting the target variable, which leads to a significant impact on TabSTAR performance (see §6). Each verbalization element is encoded independently into a semantic representation, with attention applied between tokens within each sequence element. In parallel, we encode standardized numerical values by projecting them into the same dimension using a small MLP. For the patient in Table 2, this results in a numerical embedding for Age alongside semantic representations for each of the 5 verbalizations.
@@
To obtain a unified representation for each sequence element, we apply a fusion block consisting of a single encoder-only transformer layer. Crucially, each numerical feature is fused independently, as the block attends only to its numerical and semantic embeddings. In our running example, the representation of Age now jointly captures both its semantic context (the fact that the value represents age) as well as its numerical value (the patient’s age, 45, or 0.27 after standardization).
@@
The fused, semantically-rich and numerically-grounded representations of all elements interact via a 6-layer Transformer encoder [79]. Each input element is now a token, with feature tokens and target tokens all attending to each other. Unlike standard language models, which integrate positional encoding, the Interaction module’s inputs are order-invariant, a desideratum for TFMs, as defined by [77]. The encoder produces contextualized representations for each target value. In our example, this yields dedicated embeddings for the Release and Hospitalization target values. The role of these representations is to carry information about how likely each value is to be the true value.
@@ 4 Experiments
While TabSTAR could be pretrained on a massive scale, for this work we limit ourselves to a modest pretraining corpus focusing on classification, as we believe that TabSTAR’s inductive biases are best suited to shine in this task. We manually curate a pretraining corpus of 350 high-quality tabular datasets (253 classification, 97 regression), in a tedious process in which we uncover numerous duplications in the most popular tabular repositories, OpenML [78] and Kaggle, [^10] https://www.kaggle.com/datasets as elaborated by [76]. We begin by sourcing datasets from popular benchmarks [25, 47, 23, 22, 31, 63, 55, 32], but observe that the presence of textual tabular datasets in them is very rare. Thus, we furthermore expand our corpus, focusing on classification datasets with rich semantic content. See Appendix C for more details.
@@
Tabular datasets with free-text have seen little prior research, and accordingly, benchmarks are rare. To address this, we compile all available datasets from three sources: (1) the AutoML Multimodal Benchmark [63], (2) the CARTE paper [47], and (3) the analysis of free-text and high-cardinality features by [32]. After deduplication, our final benchmark includes 50 datasets. Despite its breadth, this collection has two key limitations: First, the benchmark is heavily skewed towards regression tasks, with 36 datasets. Secondly, 29 out of these 36 datasets were solely contributed by the CARTE benchmark, which focuses more heavily on high-cardinality features, rather than longer texts, as it was pretrained over knowledge graphs. While our main motivation is classification tasks with textual features, we decide nevertheless to evaluate on the full set of 50 datasets, although it is biased toward regression problems and high-cardinality features (see Appendix D).
@@
We compare TabSTAR against GBDTs: Random Forest [12], LightGBM [46], XGBoost [15], and CatBoost [60]; MLPs: RealMLP [39] and TabM; and TFMs: [29], CARTE [47], TabDPT [54], TabPFN-v2 [38], and TabICL [61] which only supports classification tasks. For several models [^11] All MLPs and GBDTs, except the naive Random Forest baseline, used as a weak baseline. we consider both default variants as well as tuned ones, where hyperparameters are optimized separately for each task using random search with 5-fold cross-validation under a 4-hour budget on 8 CPU cores. Since the public TabPFN-v2 model does not support text, we use their closed-sourced API client. [^12] https://github.com/PriorLabs/tabpfn-client For models lacking native support for textual features, we embed text using e5-small-v2 [80], allowing a fair comparison. For more details about the hyperparameters for each baseline as well as exclusion of baselines due to potential leakage concerns, see Appendix E.
@@
Each of the 50 datasets in the benchmark is evaluated with 10 random train-test splits (90% training, 10% testing), resulting in 500 runs per model. While 30 of the datasets have more than 10,000 examples, most TFMs can’t effectively scale beyond it: TabPFN-v2, for example, employs ICL and thus receives as input at most 10,000 examples. While CARTE imposes no strict size cap, its tuning is slow, inefficient, and impractical for larger datasets. [^13] In their own paper, CARTE was evaluated only over up to 2,048 examples, without scaling guarantees. Because of these important limitations, we consider two experiment conditions: (1) 10K: Each model is trained [^14] While ICL-based TFMs aren’t technically trained, we adopt this term for conciseness. over at most 10,000 training examples, and (2) Unlimited: We evaluate TabSTAR and the most competitive, scalable baselines on the full version of the 30 datasets. [^15] We technically cap the amount of examples to 100,000 for computational efficiency.
@@
To maximize the value of cross-dataset learning, instead of pretraining TabSTAR once, we create five dataset splits. Each variant is pretrained on the 350 pretraining datasets and 40 of the benchmark datasets, while the other 10 serve exclusively as its test set. Crucially, the whole collection was carefully curated to prevent any data leakage from duplicate or overly similar datasets. As a result, each dataset is evaluated by finetuning the single pretrained variant, which excludes it from its pretraining. For finetuning, while dataset-specific hyperparameter tuning can boost performance, we believe that robust out-of-the-box defaults are essential for TFMs and their evaluation. Therefore, we use a default hyperparameter configuration that was found robust over a disjoint set of tabular datasets, as detailed in Appendix B.2.